OpenStudy (anonymous):

What are the values of the functions y =x - 4 for x = 0, 1, 2 and 3? -4, -1, 2, 5 0 ,3 , 6, 9 -4, 2, 5, 8 0, 1, 5, 9

OpenStudy (cwrw238):

just plug in x = 0 and x = 1 , 2 and 3 and see if they fit

OpenStudy (anonymous):

Huh?

OpenStudy (cwrw238):

check out the first of the option x == o y = o - 4 OK x = 1, y = 1 - 4 = -3 not OK so try next option etc